OpenStudy (anonymous):

to solve 3(x-6)-x=20. The first step would be to

OpenStudy (anonymous):

multiply everything inside the brackets 3x-18-x=20 subtract 3x-x=2x and 20-18=2 2x=2 devide each side by 2 x=1

OpenStudy (anonymous):

I do not know what step I would do first. Here are the possible answers add six to both sides divide both sides by 3 add x to both sides distribute the 3 to remove parenthesis

OpenStudy (anonymous):

follow the steps i gave you

OpenStudy (anonymous):

Answer is distribute the 3 to remove parenthesis